OpenStudy (anonymous):

Need help in this Operator problems. 1. D^2(D-3)^2(D+3)y = e^-3x 2. (D^2 - 1)y = 2e^x 3. (D^3 - D)y = 2x 4. (D^2 + 6D +25)y = 8e^-2x (sinx)

OpenStudy (anonymous):

oh. forgot to tell that I need to find the particular solution...

OpenStudy (unklerhaukus):

1. \[D^2(D-3)^2(D+3)y_p = e^{-3x}\\\,\\ y_p = \frac1{D^2(D-3)^2(D+3)}e^{-3x}\\\,\\\quad=\]

OpenStudy (anonymous):

but sir.. how do you do things in there? I only know D+a and subtracting it to it's exponent. But there's like multiple operator in there..

OpenStudy (anonymous):

should I add the exponents of the operators like D^3 * D^2 = D^5??? o.o

OpenStudy (unklerhaukus):

\[= \frac1{D^2}\frac1{(D-3)^2}\frac1{(D+3)}e^{-3x}\\\\=\]

OpenStudy (anonymous):

so should I go one by one?

OpenStudy (anonymous):

can I do this? \[(\int\limits e^{-3x})[3(\int\limits e^{-3x}) ]\]

OpenStudy (anonymous):

What operation is D?

OpenStudy (anonymous):

what do you mean sir? o.o

OpenStudy (anonymous):

Is D a variable, what is it?

OpenStudy (anonymous):

D is an operator which means "the derivative" of something.

OpenStudy (anonymous):

What is D+a? How do you add differentiation to a constant?

OpenStudy (anonymous):

it's under a certain topic.. you don't really add it.. it's just a given equation. o.o

OpenStudy (anonymous):

So what is the definition?

OpenStudy (anonymous):

err, Sire. I thought you should've known this operator D topic already... o.o

OpenStudy (anonymous):

The point is to get you thinking about it.

OpenStudy (anonymous):

Treat \(D\) like a variable and expand out the expression into a polynomial.

OpenStudy (anonymous):

You can use commutative, associative, even distributive properties on this operator.

OpenStudy (anonymous):

...okay. :< I ain't asking for anything big, I'm looking for someone to guide me. I really clueless when something I don't know is added. Like those equations above. I only know this things like: \[\frac{ 1 }{ (D-3)^3 } = 12x^{-3}e^{3x} 'cause there's only one equation for D, which is D-3 cube.. ._.

OpenStudy (anonymous):

This?\[ \frac 1 {(D-3)^3}=12x^{-3}e^{3x} \]

OpenStudy (anonymous):

Yes. o.o

OpenStudy (anonymous):

Okay so what do the other equations add which you don't quite get?

OpenStudy (anonymous):

those above. ._.

OpenStudy (anonymous):

Yes, but what aspect do they add which you don't get? What makes them different?

OpenStudy (anonymous):

Have you ever considered, that the fact that they're all factored polynomials means that the advantage of using D operator, is that finding the roots will give insight?

OpenStudy (anonymous):

if that's the case.. well.. I've been thinking of that too. I'm just too afraid that I might not get the real answer.

OpenStudy (anonymous):

Okay, here is the skinny. I don't know how to do it, so I learned how to do it. If you have a polynomial of differential operators \(L(D)\) Then a solution to the equation: \[ L(D)y=e^{kx} \]Would be \[ y_1=\frac{e^{kx}}{L(k)} \]

OpenStudy (anonymous):

Let's start with number one.

OpenStudy (anonymous):

We have \( e^{-3x}\) so \(k=-3\)

OpenStudy (anonymous):

\[ L(D)=D^2(D-3)^2(D+3) \]And thus: \[ L(k) = L(-3) = (-3)^2(-3-3)^2(-3+3) \]

OpenStudy (anonymous):

Since apparently \(L(k)=0\), it doesn't yeild anything.

OpenStudy (anonymous):

err.. so. it's (9)(36) ?

OpenStudy (anonymous):

oh. lol. 0. xD

OpenStudy (anonymous):

Yeah, this method can't help with this equation.

OpenStudy (anonymous):

Actually, I learned some more about this method. Actually it can still work if \(L(k)=0\)

OpenStudy (anonymous):

You have to figure out the multiplicity of the \(0\). In this case it is \(1\).

OpenStudy (anonymous):

Since it was caused by \((-3+3)^{\color{red}1}\)

OpenStudy (anonymous):

in this case the solution is: \[ \frac{x^ne^{kx}}{L^{(n)}(k)} \]Where \(n\) is the multiplicity.

OpenStudy (anonymous):

So we need the first derivative of \(L(k)\). Can you do that?

OpenStudy (anonymous):

Please tell me you know how to take the derivative.

OpenStudy (anonymous):

in what equation? o.o

OpenStudy (anonymous):

.............. really? Okay, find t he derivative of \(L(D)\) with respect to \(D\) \(\color{blue}{\text{Originally Posted by}}\) @wio \[ L(D)=D^2(D-3)^2(D+3) \]And thus: \[ L(k) = L(-3) = (-3)^2(-3-3)^2(-3+3) \] \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

so it's uhh.. 2D[2(D-3)]D then.. 2(-3)[2(-3-3)](-3) (-6)(-12)(-3) = 222 ??

OpenStudy (anonymous):

You didn't do the product rule correctly.

OpenStudy (anonymous):

It would be easier to just multiply the whole thing out first.

OpenStudy (anonymous):

(D−3)2⋅D^2+2⋅(D−3)⋅D^2⋅(D+3)+2⋅(D−3)^2⋅D⋅(D+3) err.. 0.

OpenStudy (anonymous):

Expand it out into a polynomial first, then differentiation. Let's avoid the product rule mess.

OpenStudy (anonymous):

okay..

OpenStudy (anonymous):

\[ D^2(D-3)^2(D+3)= 27 D^2-9 D^3-3 D^4+D^5 \]

OpenStudy (anonymous):

Find the derivative of this polynomial, then put in \(-3\)

OpenStudy (anonymous):

-360? o.o

OpenStudy (anonymous):

432..

OpenStudy (anonymous):

I'm getting -108

OpenStudy (anonymous):

I used wolfram to do the derivative

OpenStudy (anonymous):

so I'd say our solution is: \[ -\frac{xe^{-3x}}{108} \]

OpenStudy (anonymous):

okay..

OpenStudy (anonymous):

With the second one, I think the same logic applies.

OpenStudy (anonymous):

Well, if you showed me your work I could tell you what is wrong.

OpenStudy (anonymous):

well.. I still don't get #1. I tried solving #2.

OpenStudy (anonymous):

Here's what I get: \[(D^2-1)y = 2e^x \] factoring out D^2-1 \[(D-1)(D+1)y = 2e^x\] now, \[(D+1)yp = 2e^x\] I'm just setting aside D-1, deriving 2e^x.. \[(D-1)yp = 2e^x +2e^x\] then, \[y _p=2e^x+2e^x+2e^x-2e^x\] \[y_p = 4e^x\] but from the book it should be xe^x.. dunno why.

OpenStudy (anonymous):

\[ (D^2 - 1)y = 2e^x \]So \[ \frac 12 (D^2 - 1)y = e^x \]

OpenStudy (anonymous):

Now factoring gets \((D-1)(D+1)\) so the mutiplicity of the root is 1.

OpenStudy (anonymous):

Doing the derivative of \[ \frac 12 (D^2 - 1) \]returns \(D\)

OpenStudy (anonymous):

So our solution is \[ \frac{xe^x}{1} \]

OpenStudy (anonymous):

err.. how did it became D?

OpenStudy (anonymous):

oh. wait. I get it already.

OpenStudy (anonymous):

so how did it became xe^x?

OpenStudy (anonymous):

Remember it is \(x^n\) where \(n\) is the multiplicity of the root? Well \(n=1\)

OpenStudy (anonymous):

how can I show it properly? o.o I can only get.. uhh..\[\frac{ D^2-1 }{ 2 }\]

OpenStudy (anonymous):

Okay so what is \[ L'(D) \]In this case?

OpenStudy (anonymous):

well.. D is an operator.. I can't derive it properly.. but its on the terms of x, the answer is x

OpenStudy (anonymous):

when it becomes just D, it becomes the derivative of e^x which is e^x.. o.o

OpenStudy (anonymous):

oh. here. look. \[\frac{ 1 }{ 2 }(D-1)(D+1)y = e^x\]

OpenStudy (anonymous):

well, how do we really get the x from there? o.o

OpenStudy (anonymous):

why did D became x? >.>

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

Sorry

OpenStudy (anonymous):

D didn't become x

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @wio in this case the solution is: \[ \frac{x^ne^{kx}}{L^{(n)}(k)} \]Where \(n\) is the multiplicity. \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

Since \(n=1\) we have \(xe^x\)

OpenStudy (anonymous):

We also have the first derivative of L

OpenStudy (anonymous):

what's the name of that solution?

OpenStudy (anonymous):

Exponential input theorem

OpenStudy (anonymous):

oh..

OpenStudy (unklerhaukus):

\[ \frac1{D^2}\frac1{(D-3)^2}\frac1{(D+3)}\cdot e^{-3x}\] the inner most bit first ; \[\frac1{(D+3)}\cdot e^{-3x}=e^{-3x}\frac{1}{(D-3)+3}\cdot1\\ \qquad\qquad\qquad=e^{-3x}\frac1D\\ \qquad\qquad\qquad=e^{-3x}\int dx\] so \[= \frac1{D^2}\frac1{(D-3)^2}e^{-3x}\]

OpenStudy (unklerhaukus):

note that the particular solution dosent need a consant of integration